the sum of ages of Gandhi and Ahmed is 38. Seven years ago, Ghandhi was three times as old as Ahmed. Find there present ages.
Answer by chirag17oct@gmail.com(1)

let gandhi age be x & ahmed age be y
now listen bro or sis,
x+y=38(given)
x=38-y
acco to Q- 7 yrs ago
(x-7)& (y-7)
now, gandhi was 3 tims as old as ahmad
so,3(x-7)=y-7
3x-21=y-7(put the value of x)
3(38-y)-21=y-7
y=25
x+y=38
x=38-25
13ans.
so gandhi age =13yrs & ahmad age=25yrs